# Slaying Axion-Like Particles via Gravitational Waves and Primordial Black Holes from Supercooled Phase Transition

### v1

@@ 5.2 The astrophysical foreground
As mentioned in the Introduction, the cosmological stochastic GW signal discussed here is subject to an astrophysical foreground from unresolved compact binary coalescences [211]. LIGO/VIRGO has already observed GW events involving binary black hole (BH-BH) and binary neutron stars (NS-NS) mergers [212, 213]; hence, an astrophysical foreground is guaranteed to exist. The sum of the diffuse astrophysical foreground is shown in Fig. 5 as the gray shaded region in the bottom half plane (see also Refs. [37, 128]). Different subtraction methods have been proposed to deal with this foreground at future GW detectors and to extract any potential signal of primordial origin [214, 215, 216, 217, 218, 219, 220, 221, 222]. We expect that the NS and BH foreground can be subtracted from the sensitivities of BBO and ET or CE in the ranges $\Omega_{\rm GW}\simeq 10^{-15}$ [214] and $\Omega_{\rm GW}\simeq 10^{-13}$ [216]. However in the LISA window, the galactic and extra-galactic binary white dwarf (WD-WD) foregrounds may dominate over the NS-NS and BH-BH foregrounds [223, 224, 225] and should be subtracted [226] from the expected sensitivity $\Omega_{\rm GW}\simeq 10^{-13}$ to be reached at LISA [227, 228]. Given that such subtractions can be made possible with a network of GW detectors due to the fact that the GW spectrum generated by the astrophysical foreground increases with frequency as $f^{2/3}$ [229, 215], which is different from the GW spectrum generated by nucleating bubbles during strong FOPT ($\Omega_{\rm GW}\sim f^{2.3}$ and $\Omega_{\rm GW}\sim f^{-2.4}$ for frequencies below and above the peak, respectively; see Fig. 5), we hope to be able to eventually pin down the GW signals from supercooled PT.
@@ 5.3 Fitting the NANOGrav signal
Figure 6: The 15-year data from the NANOGrav collaboration (blue violins) and the best-fit GW spectrum (solid orange, with the $2\sigma$ band shown by the orange shaded region) from the supercooled FOPT in our ALP model. For comparison, the expected astrophysical contribution from SMBH is shown by the gray band. Other labels are the same as in Fig. 5.
In Fig. 6, we focus on the nHz frequency regime and the 15-year NANOGrav data on the SGWB [23]. Our ALP model can explain this signal for $f_{a}\simeq(10~{}{\rm GeV}-1~{}{\rm TeV})$ and $g\simeq 0.56-0.60$. The best-fit GW spectrum is shown by the solid orange curve and the $2\sigma$ band is shown by the dashed orange curves. For comparison, we also show the expected band of spectra from astrophysical sources, namely, supermassive black hole (SMBH) binaries [230]. Our cosmological signal tends to fit the NANOGrav data better, but a more accurate modeling of the SMBH background could alter this result. Our $2\sigma$ best-fit spectrum is also consistent with the $\Delta N_{\rm eff}$ bound from BBN and Planck data (horizontal gray shaded region). However, as shown in Fig. 3, the required values of $f_{a}$ and $g$ in our model to fit the NANOGrav signal lead to PBH overproduction, and thus, disfavored in the simplest version.
@@ 5.4 PBH constraints
In Fig. 7 we depict the observational constraints on $f_{\rm PBH}$ (see Ref. [105] for a review). PBH evaporation via Hawking radiation give constraints in (see also Refs. [231, 232, 233]): CMB [234], EDGES [235], INTEGRAL [236, 237], Voyager [238], 511 keV [239], and EGRB [240]. The microlensing-related constraints from HSC (Hyper-Supreme Camera) [241], EROS [242] and Icarus [243], as well as the hint of PBH from OGLE [244] are also shown. We also show the future micro-lensing sensitivity reach of Nancy Grace Roman Space Telescope (NGRST) [245]. Constraints coming from modification of the CMB spectrum due to PBH accretion are shown on the right [246] (see also Ref. [247]). Finally the mass range around $M_{\odot}$ is constrained by LIGO-VIRGO observations on PBH-PBH merger  [248, 249, 120, 250, 119, 251, 121], while future GW interferometers like ET will be able to set better limits on the PBH abundance [252, 253, 254, 255, 256, 257] as depicted by the dashed gray curve in the plot.
@@
Given the PBH constraints in Fig. 7, we take five benchmark points A, B, C, D, E in the allowed region, with the masses of the PBHs formed during the supercooling as $10^{-17.5}M_{\odot}$, $10^{-16}M_{\odot}$, $10^{-13}M_{\odot}$, $10^{-10.5}M_{\odot}$ and $10^{-8}M_{\odot}$ respectively. While B, C and D correspond to the PBH accounting for the entire DM content of the Universe, for benchmark points A and E, PBHs only comprise $10^{-5}$ of the DM relic density of the Universe. The GW spectra for these benchmark points are shown in Fig. 5.
@@ 6 Complementarity with laboratory and astrophysical ALP searches
Finally, in Fig. 10 we show the constraints on the ALP-nucleon coupling. In this case, the existing haloscope constraints (blue shaded), coming from experiments like CASPEr [289], K-${}^{3}$e comagnetometer [290], nEDM [291], NASDUCK [292, 293], JEDI [294] and ChangE [295, 296], and astrophysical constraints from isolated neutron star cooling [297] are relevant. But the PBH constraints from BBN and CMB turn out to be the strongest, even cutting into the future laboratory sensitivities.
@@ 7 Discussion and Conclusion
Furthermore, in our model there is a one-to-one correspondence between the ALP decay constant $f_{a}$ and the PBH mass $M_{\rm PBH}$. This leads to a three-prong complementarity between the GW signal, PBH signals and other laboratory/astrophysical ALP probes. In particular, depending on the ALP couplings to the SM particles, the ALP constraints can be translated into a lower bound on the PBH mass, independent of $f_{\rm PBH}$ (see Fig. 7). Analogously, we can use the existing PBH constraints to ‘slay’ additional ALP parameter space, which for $f_{\rm PBH}=1$ turn out to be even more stringent than the future laboratory searches (see Figs. 8, 9 and 10). This $f_{\rm PBH}=1$ region can be tested via the GW signal in future GW detectors like LISA, ET and CE (see Fig. 4). We also discussed the recent NANOGrav detection of a SGWB and its possible interpretation in terms of the FOPT in our ALP scenario (see Fig. 6); however, we find that the NANOGrav-preferred region leads to PBH overclosure.
@@ Acknowledgments
We thank Priyamvada Natarajan and Kai Schmitz for comments on the draft. A.C. thanks the Galileo Galilei Institute for Theoretical Physics for the hospitality and the INFN for partial support during the completion of this work. This work has been partly funded by the European Union – Next Generation EU through the research grant number P2022Z4P4B “SOPHYA - Sustainable Optimised PHYsics Algorithms: fundamental physics to build an advanced society” under the program PRIN 2022 PNRR of the Italian Ministero dell’Università e Ricerca (MUR) and partially supported by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing. The work of B.D. was partly supported by the U.S. Department of Energy under grant No. DE-SC 0017987.

### v2

@@ 5.2 The astrophysical foreground
As mentioned in the Introduction, the cosmological stochastic GW signal discussed here is subject to an astrophysical foreground from unresolved compact binary coalescences [217]. LIGO/VIRGO has already observed GW events involving binary black hole (BH-BH) and binary neutron stars (NS-NS) mergers [218, 219]; hence, an astrophysical foreground is guaranteed to exist. The sum of the diffuse astrophysical foreground is shown in Fig. 5 as the gray shaded region in the bottom half plane (see also Refs. [37, 128]). Different subtraction methods have been proposed to deal with this foreground at future GW detectors and to extract any potential signal of primordial origin [220, 221, 222, 223, 224, 225, 226, 227, 228]. We expect that the NS and BH foreground can be subtracted from the sensitivities of BBO and ET or CE in the ranges $\Omega_{\rm GW}\simeq 10^{-15}$ [220] and $\Omega_{\rm GW}\simeq 10^{-13}$ [222]. However in the LISA window, the galactic and extra-galactic binary white dwarf (WD-WD) foregrounds may dominate over the NS-NS and BH-BH foregrounds [229, 230, 231] and should be subtracted [232] from the expected sensitivity $\Omega_{\rm GW}\simeq 10^{-13}$ to be reached at LISA [233, 234]. Given that such subtractions can be made possible with a network of GW detectors due to the fact that the GW spectrum generated by the astrophysical foreground increases with frequency as $f^{2/3}$ [235, 221], which is different from the GW spectrum generated by nucleating bubbles during strong FOPT ($\Omega_{\rm GW}\sim f^{3}$ and $\Omega_{\rm GW}\sim f^{-2.4}$ for frequencies below and above the peak, respectively; see Fig. 5), we hope to be able to eventually pin down the GW signals from supercooled PT.
@@ 5.3 Fitting the NANOGrav signal
Figure 6: The 15-year data from the NANOGrav collaboration (blue violins) and the $2\sigma$ bands (shown by the orange and red shaded regions) corresponding to the best-fit GW spectrum from the supercooled FOPT in our ALP model. For comparison, the expected astrophysical contribution from SMBH is shown by the two gray bands. Other labels are the same as in Fig. 5.
In Fig. 6, we focus on the nHz frequency regime and the 15-year NANOGrav data on the SGWB [23]. Our ALP model can explain this signal for $f_{a}\simeq(10~{\rm GeV}-1~{\rm TeV})$ and $g\simeq 0.56-0.60$. As explained in the context of Fig. 3, we calculate the $2\sigma$ band that fits the NANOGrav signal in two ways, i.e. with only the first 14 bins (red shaded) and including all 30 bins (orange shaded). For comparison, we also show the expected band of spectra from astrophysical sources, namely, supermassive black hole (SMBH) binaries as evaluated in [236] and, more recently, in [27]. Our cosmological signal tends to fit the NANOGrav data slightly better, especially if we take all 30 bins. Our $2\sigma$ best-fit spectra are also consistent with the $\Delta N_{\rm eff}$ bound from BBN and Planck data (horizontal gray shaded region), although the NANOGrav data in the high-frequency bins are in some tension with the BBN and Planck constraints.
@@ 5.4 PBH constraints
In Fig. 7 we depict the observational constraints on $f_{\rm PBH}$ (see Ref. [105] for a review). PBH evaporation via Hawking radiation give constraints in (see also Refs. [237, 238, 239]): CMB [240], EDGES [241], INTEGRAL [242, 243], Voyager [244], 511 keV [245, 246], and EGRB [247]. The microlensing-related constraints from HSC (Hyper-Supreme Camera) [248], EROS [249] and Icarus [250], as well as the hint of PBH from OGLE [251] are also shown. We also show the future micro-lensing sensitivity reach of Nancy Grace Roman Space Telescope (NGRST) [252]. Constraints coming from modification of the CMB spectrum due to PBH accretion are shown on the right [253] (see also Ref. [254]). Finally the mass range around $M_{\odot}$ is constrained by LIGO-VIRGO observations on PBH-PBH merger  [255, 256, 120, 257, 119, 258, 121], while future GW interferometers like ET will be able to set better limits on the PBH abundance [259, 260, 261, 262, 263, 264] as depicted by the dashed gray curve in the plot.
@@
Given the PBH constraints in Fig. 7, we take five benchmark points A, B, C, D, E in the allowed region, with the masses of the PBHs formed during the supercooling as $10^{-17.5}M_{\odot}$, $10^{-16}M_{\odot}$, $^{-13}M_{\odot}, 1010$$^{-10.5}M_{\odot}$ and $10^{-8}M_{\odot}$ respectively. While B, C and D correspond to the PBH accounting for the entire DM content of the Universe, for benchmark points A and E, PBHs only comprise $10^{-5}$ of the DM relic density of the Universe. The GW spectra for these benchmark points are shown in Fig. 5.
@@
For axion decay constant values $f_{a}=10^{5}-10^{8}$ GeV, PBH can be the entire DM candidate of the universe (see Fig. 7) while for other values, PBH can still be the fractional DM candidate while the dominant component of the DM may come from ALP relic produced due to vacuum misalignment mechanism [265], see Fig. 8.
@@ 6 Complementarity with laboratory and astrophysical ALP searches
Finally, in Fig. 10 we show the constraints on the ALP-nucleon coupling. In this case, the existing haloscope constraints (blue shaded), coming from experiments like CASPEr [297], K-3He comagnetometer [298], nEDM [299], NASDUCK [300, 301], JEDI [302] and ChangE [303, 304], and astrophysical constraints from isolated neutron star cooling [305] are relevant. But the PBH constraints from BBN and CMB turn out to be the strongest, even cutting into the future laboratory sensitivities.
@@ 7 Discussion and Conclusion
Furthermore, in our model there is a one-to-one correspondence between the ALP decay constant $f_{a}$ and the PBH mass $M_{\rm PBH}$. This leads to a three-prong complementarity between the GW signal, PBH signals and other laboratory/astrophysical ALP probes. In particular, depending on the ALP couplings to the SM particles, the ALP constraints can be translated into a lower bound on the PBH mass, independent of $f_{\rm PBH}$ (see Fig. 7). Analogously, we can use the existing PBH constraints to ‘slay’ additional ALP parameter space, which for $f_{\rm PBH}=1$ turn out to be even more stringent than the future laboratory searches (see Figs. 8, 9 and 10). This $f_{\rm PBH}=1$ region can be tested via the GW signal in future GW detectors like LISA, ET and CE (see Fig. 4). We also discussed the recent NANOGrav detection of a SGWB and its possible interpretation in terms of the FOPT in our ALP scenario (see Fig. 6).
@@ Acknowledgments
We thank Yann Gouttenoire, Kayhan Gultekin, Marek Lewicki, Priyamvada Natarajan and Kai Schmitz for comments on the draft. A.C. thanks the Galileo Galilei Institute for Theoretical Physics for the hospitality and the INFN for partial support during the completion of this work. This work has been partly funded by the European Union – Next Generation EU through the research grant number P2022Z4P4B “SOPHYA - Sustainable Optimised PHYsics Algorithms: fundamental physics to build an advanced society” under the program PRIN 2022 PNRR of the Italian Ministero dell’Università e Ricerca (MUR) and partially supported by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing. The work of B.D. was partly supported by the U.S. Department of Energy under grant No. DE-SC 0017987.